Saturday, April 19, 2014

Night Log 2014-04-18

SA: Paul
SO: Veronica
Others: Luke, Lisa, Eddy


- we had thin cloud until 02:00 which made target acquisition difficult with the bright moon. attempted a couple of targets, but could not successfully acquire them.
- did pointing engineering observations when we had gaps in the queue.
- the seeing was good, mostly below 1.5''
- reconstruction of the HRS thermal enclosure was completed tonight & the stulz aircon was switched on just after midnight to begin stabilizing the temperature of the HRS inner room

- obtained data for the following programs:
- 2013-2-RU-001 (supernova follow-up observations) P3
- 2013-2-RSA_OTH-003

No comments:

Post a Comment